OpenStudy (anonymous):

-3(2x - 5) - 4(x + 7) simplify.

OpenStudy (anonymous):

@amistre64 do i do the same thing as my previous problem?

OpenStudy (luigi0210):

Distribute the numbers and combine like terms

OpenStudy (amistre64):

yes, but this time you have something other than a 1 to distribute

OpenStudy (amistre64):

distribution is just a way of speeding up addition ... its multiplication. 4(a+b) = (a+b) + (a+b) + (a+b) + (a+b) = a+b+a+b+a+b+a+b = a+a+a+a + b+b+b+b = 4a + 4b

OpenStudy (anonymous):

oh i see. so i multiply everything in the parentheses by the number outside and before it?

OpenStudy (amistre64):

thats a quick version yes :)

OpenStudy (anonymous):

-6x+15 -4x-28

OpenStudy (anonymous):

then i subtract it like the other problem right?

OpenStudy (amistre64):

-3(2x - 5) -4(x + 7) -6x + 15 -4x - 28 yep thats perfect, now combine like terms

OpenStudy (anonymous):

-10x-13

OpenStudy (amistre64):

i like it

OpenStudy (anonymous):

Great! thanks for your help again! :)

OpenStudy (amistre64):

youre welcome .. your doing great :)

OpenStudy (anonymous):

-10x-13